4. Determine the value of the equilibrium constant, Q for the chemical reaction given below, if 0.109 moles of Al, 0.076 moles of Th, 0.0062 of At, and 0.0090 H2O in water vapor are in a 10 liters container and at an equilibrium.


Al= 0.109


Th= 0.076


At= 0.0062


H2O water vapor= 0.0090


10 liters


substitute the value of the compounds


1
Expert's answer
2021-04-05T07:57:02-0400

"Moles=0.109+0.076+0.0062+0.0090=0.2002"


"Q constant = \frac{0.2002}{4}×10=0.5005"